# Proximality and selflessness for group C*-algebras

### v6

@@ 1. Introduction
We say a sequence $(z_{n})_{n}$ in $\Gamma$ is axial if there is a topologically-free extreme boundary $\Gamma\curvearrowright X$ with distinct points $z_{\pm}\in X$ that satisfies the following two conditions. (1): For every neighborhoods $U_{\pm}$ of $z_{\pm}$ one has $z_{n}(X\setminus U_{-})\subset U_{+}$ eventually, or equivalently that $z_{n}^{-1}(X\setminus U_{+})\subset U_{-}$ eventually. (2): The $\Gamma$-action on $\{z_{\pm}\}$ is free in the sense that $g\{z_{\pm}\}\cap\{z_{\pm}\}\neq\emptyset$ implies $g=1$. Note that an extreme boundary $X$ with $|X|>2$ accommodates an axial sequence if and only if it is topologically-free.
@@ Acknowledgments
A part of this work was carried out while the author was visiting the Mathematisches Forschungsinstitut Oberwolfach for the workshop “C*-Algebras”, August 3–8, 2025 and the Isaac Newton Institute for the program “Operators, Graphs, Groups” in October 2025. He acknowledges the kind hospitality and the exciting environment provided by the institutes. He thanks G. Patchell for the stimulating talk at MFO and a useful comment on this paper and E. Breuillard and I. Vigdorovich for insights into $\mathrm{PSL}(d,\mathbb{R})$.
@@ 8. The Powers property and selflessness
We say that a group $\Gamma$ has property $\mathrm{P}_{\mathrm{PHP}}$ if for every finite subset $F\subset\Gamma$, $\varepsilon>0$, and $n\in\mathbb{N}$, there are $t_{i}\in\Gamma$ and $C_{i}\subset D_{i}\subset\Gamma$, $i=1,\ldots,n$, satisfying that the members of
@@
Given a finite subset $F\subset\Gamma$, $\varepsilon>0$, and $n\in\mathbb{N}$, we take $t_{i}$ and $C_{i}\subset D_{i}\subset\Gamma$ as in Definition. Let $P_{i}\leq Q_{i}$ denote the diagonal projections in $\mathbb{B}(\ell_{2}\Gamma)$ corresponding to the subsets $C_{i}\subset D_{i}$ and set $P_{i}^{\prime}\coloneq t_{i}^{-1}Q_{i}^{\perp}t_{i}$ and $Q_{i}^{\prime}\coloneq t_{i}^{-1}P_{i}^{\perp}t_{i}$. Note that $\{P_{i}\}\cup\{P^{\prime}_{j}\}$ are mutually orthogonal. Here $\Gamma$ is represented in $\mathbb{B}(\ell_{2}\Gamma)$ by the left regular representation. We denote by $\omega$ the vector state on $\mathbb{B}(\ell_{2}\Gamma)$ corresponding to $\delta_{1}$, which satisfies $\omega|_{\mathrm{C}^{*}_{\lambda}(\Gamma)}=\tau$. We define $T\in\mathbb{B}(\ell_{2}\Gamma)$ by
@@
$\|\sum_{i}(Q_{i}-P_{i})t_{i}\|$ $=\|\left[\begin{matrix}Q_{1}-P_{1}&\cdots&Q_{n}-P_{n}\end{matrix}\right]\left[\begin{matrix}t_{1}&&\\ &\ddots&\\ &&t_{n}\end{matrix}\right]\left[\begin{matrix}Q^{\prime}_{1}\\ \vdots\\ Q^{\prime}_{n}\end{matrix}\right]\|$
$\leq\|\sum_{i}Q_{i}-P_{i}\|^{1/2}\|\sum_{i}Q^{\prime}_{i}\|^{1/2}\leq\varepsilon n^{1/2}$
@@
We sketch the proof by adopting that of Lemma 4 in [dlH]. Let a finite subset $F\subset\Gamma$ and $n$ be given. Then, there are $t_{1},\ldots,t_{n}$ in $\Gamma$ and mutually disjoint open subsets $U_{i}^{\pm}$ such that $t_{i}(L\setminus U_{i}^{-})\subset U_{i}^{+}$ for every $i$. Moreover, we may assume that $a(U_{i}^{+}\cup U_{i}^{-})\cap(U_{j}^{+}\cup U_{j}^{-})=\emptyset$ for every $a\in F\setminus\{1\}$ and every $i,j$. We fix a point $x_{0}\in L$ and set $C_{i}\coloneq D_{i}\coloneq\{s\in\Gamma:sx_{0}\in U_{i}^{+}\}$. Then they witness $\mathrm{P}_{\mathrm{PHP}}$ with $\varepsilon=0$. ∎
@@
Let a finite subset $F\subset\Gamma$ be given. We claim that for every $n$ there are $s_{1},\ldots,s_{n}\in\Gamma$ such that the members of
@@
are mutually disjoint and that every $I\subset\{1,\ldots,n\}$ with cardinality $d$ satisfies $\bigcap_{i\in I}V_{i}=\emptyset$ and $\bigcap_{i\in I}V_{i}^{\prime}=\emptyset$. Then by compactness, for each $i$ there is $n_{i}\in\mathbb{N}$ large enough so that $t_{i}\coloneq s_{i}t_{0}^{n_{i}}s_{i}^{-1}$ satisfies $t_{i}^{-1}(B\setminus U_{i})\subset V_{i}^{\prime}$ and $t_{i}^{-1}(B\setminus V_{i})\subset U_{i}^{\prime}$. We further choose $x_{0}\in B$ and set $C_{i}\coloneq\{g\in\Gamma:gx_{0}\in U_{i}\}$ and $D_{i}\coloneq\{g\in\Gamma:gx_{0}\in V_{i}\}$. Thus $t_{1},\ldots,t_{n}$ and $C_{i}\subset D_{i}$ verify $\mathrm{P}_{\mathrm{PHP}}$ with
$\sup_{x\in\Gamma}|\{i:x\in D_{i}\}\cup\{j:x\in t_{j}^{-1}(\Gamma\setminus C_{j})\}|\leq 2(d-1).$

### v7

@@ 1. Introduction
We say a sequence $(z_{n})_{n}$ in $\Gamma$ is axial if there is a topologically-free extreme boundary $\Gamma\curvearrowright X$ with distinct points $z_{\pm}\in X$ that satisfies the following two conditions. (1): For every neighborhoods $U_{\pm}$ of $z_{\pm}$ one has $z_{n}(X\setminus U_{-})\subset U_{+}$ eventually, or equivalently that $z_{n}^{-1}(X\setminus U_{+})\subset U_{-}$ eventually. (2): The $\Gamma$-action on $\{z_{\pm}\}$ is free in the sense that $g\{z_{\pm}\}\cap\{z_{\pm}\}\neq\emptyset$ implies $g=1$. Note that a second countable extreme boundary $X$ with $|X|>2$ accommodates an axial sequence if and only if it is topologically-free.
@@ Acknowledgments
A part of this work was carried out while the author was visiting the Mathematisches Forschungsinstitut Oberwolfach for the workshop “C*-Algebras”, August 3–8, 2025 and the Isaac Newton Institute for the program “Operators, Graphs, Groups” in October 2025. He acknowledges the kind hospitality and the exciting environment provided by the institutes. He thanks G. Patchell for the stimulating talk at MFO and a useful comment on this paper, E. Breuillard and I. Vigdorovich for insights into $\mathrm{PSL}(d,\mathbb{R})$, and L. Robert for a comment.
@@ 8. The Powers property and selflessness
We say that a group $\Gamma$ has property $\mathrm{P}_{\mathrm{PHP}}$ if for every finite subset $F\subset\Gamma$ and $\varepsilon>0$, there are $n\in\mathbb{N}$ and $t_{i}\in\Gamma$ and $C_{i}\subset D_{i}\subset\Gamma$ for $1\leq i\leq n$ satisfying that the members of
@@
Given a finite subset $F\subset\Gamma$ and $\varepsilon>0$, we take $t_{i}$ and $C_{i}\subset D_{i}\subset\Gamma$ for $1\leq i\leq n$ as in Definition. Let $P_{i}\leq Q_{i}$ denote the diagonal projections in $\mathbb{B}(\ell_{2}\Gamma)$ corresponding to the subsets $C_{i}\subset D_{i}$ and set $P_{i}^{\prime}\coloneq t_{i}^{-1}Q_{i}^{\perp}t_{i}$ and $Q_{i}^{\prime}\coloneq t_{i}^{-1}P_{i}^{\perp}t_{i}$. Note that $\{P_{i}\}\cup\{P^{\prime}_{j}\}$ are mutually orthogonal. Here $\Gamma$ is represented in $\mathbb{B}(\ell_{2}\Gamma)$ by the left regular representation. We denote by $\omega$ the vector state on $\mathbb{B}(\ell_{2}\Gamma)$ corresponding to $\delta_{1}$, which satisfies $\omega|_{\mathrm{C}^{*}_{\lambda}(\Gamma)}=\tau$. We define $T\in\mathbb{B}(\ell_{2}\Gamma)$ by
@@
$\|\sum_{i}(Q_{i}-P_{i})t_{i}\|$ $=\|\sum_{i}(Q_{i}-P_{i})t_{i}(Q_{i}^{\prime}-P_{i}^{\prime})\|\leq\|\sum_{i}Q_{i}\|^{1/2}\|\sum_{i}Q^{\prime}_{i}\|^{1/2}\leq\varepsilon n^{1/2}$
@@
We sketch the proof by adopting that of Lemma 4 in [dlH]. Let a finite subset $F\subset\Gamma$ and $\varepsilon>0$ be given. Take $n\geq\varepsilon^{-2}$. Then, there are $t_{1},\ldots,t_{n}$ in $\Gamma$ and mutually disjoint open subsets $U_{i}^{\pm}$ such that $t_{i}(L\setminus U_{i}^{-})\subset U_{i}^{+}$ for every $i$. Moreover, we may assume that $a(U_{i}^{+}\cup U_{i}^{-})\cap(U_{j}^{+}\cup U_{j}^{-})=\emptyset$ for every $a\in F\setminus\{1\}$ and every $i,j$. We fix a point $x_{0}\in L$ and set $C_{i}\coloneq D_{i}\coloneq\{s\in\Gamma:sx_{0}\in U_{i}^{+}\}$. They witness $\mathrm{P}_{\mathrm{PHP}}$. ∎
@@
Let a finite subset $F\subset\Gamma$ and $\varepsilon>0$ be given. We claim that for every $n$ there are $s_{1},\ldots,s_{n}\in\Gamma$ such that the members of
@@
are mutually disjoint and that every $I\subset\{1,\ldots,n\}$ with cardinality $d$ satisfies $\bigcap_{i\in I}V_{i}=\emptyset$ and $\bigcap_{i\in I}V_{i}^{\prime}=\emptyset$. Then by compactness, for each $i$ there is $n_{i}\in\mathbb{N}$ large enough so that $t_{i}\coloneq s_{i}t_{0}^{n_{i}}s_{i}^{-1}$ satisfies $t_{i}^{-1}(B\setminus U_{i})\subset V_{i}^{\prime}$ and $t_{i}^{-1}(B\setminus V_{i})\subset U_{i}^{\prime}$. We further choose $x_{0}\in B$ and set $C_{i}\coloneq\{g\in\Gamma:gx_{0}\in U_{i}\}$ and $D_{i}\coloneq\{g\in\Gamma:gx_{0}\in V_{i}\}$. Thus, for $n$ large enough, $t_{1},\ldots,t_{n}$ and $C_{i}\subset D_{i}$ verify $\mathrm{P}_{\mathrm{PHP}}$ with
$\sup_{x\in\Gamma}|\{i:x\in D_{i}\}\cup\{j:x\in t_{j}^{-1}(\Gamma\setminus C_{j})\}|\leq 2(d-1)\leq\varepsilon n^{1/2}.$